(2/-7)+(2/3)
solve fraction in simplist form

Found 2 solutions by Alan3354, unlockmath:
Answer by Alan3354(69443)

(2/-7)+(2/3)
= (2/3) - (2/7)
= (14/21) - 6/21
= 8/21

Answer by unlockmath(1688)

Hello,
First we need a lowest common denominator for these: (2/-7)+(2/3)
It'll be 21
So:
-6/21+ 14/21 This comes out to:
8/21
There you go.
RJ Check out the new book I wrote at:
www.math-unlock.com